Archer Contracting repaved 50 miles of two-lane county roadway with a crew of six employees. This crew worked 8 days and used $7000 worth of paving material. Nearby, Bronson Contracting repaved 35 miles of four-lane interstate roadway working 10 days with a crew of five employees using $9000 of paving material. Both Archer Contracting and Bronson Contracting rented the same paving equipment to complete their respective assignments, costing each $1500 a day, and labor cost per employee is $160 a day for both companies. Which company is more productive and by how much?

a. Bronson Contracting is more productive by about 16.7%

b. Bronson Contracting is more productive by about 20.9%

c. Archer Contracting is more productive by about 16.7%

d. Archer Contracting is more productive by about 20.9%

e.Archer Contracting is more productive by about 71.3%
Answer #1

Archer Contracting :

Length of roadway repaved = 50 miles of two lane country road

Cost of labor = $ 160 / employee-day x 6 employees x 8 days = $7680

Cost of equipment = $1500/day x 8 days = $12000

Cost of material = $7000

Total cost of labor , equipment and material = $7680 + $12000 + $7000 = $26680

Therefore , Multifactor productivity = Length of 2 lane roadway repaved/ Total cost = 50 / 26680 = 0.00187

Boston consulting :

Length of roadway repaved = 35 miles of 4 lane country road = 70 miles of 2 lane country road

Cost of labor = $ 160 / employee-day x 5 employees x 10 days = $8000

Cost of equipment = $1500/day x 10 days = $ 15000

Material cost = $9000

Total cost of labor equipment and material = $8000 + $15000 + $9000 = $32,000

Therefore , Multifactor productivity = Length of 2 way roadway repaved/ Total cost = 70 / 32000 = 0.00218

Thus,

Multifactor productivity of Bronson contracting ( 0.00218 ) > Multifactor productivity of Archer Contracting ( 0.00187)

Hence, BRONSON CONTRACTING IS MORE PRODUCTIVE

Percentage by which Bronson Contracting is more productive = ( 0.00218 – 0.00187)/0.00187 x 100 = 16.57% ( this is nearest to the value of 16.7% as mentioned in possible answer )

Therefore correct answer will be “Bronson Contracting is more productive by 16.7%”

ANSWER : a ) BRONSON CONTRACTING IS MORE PRODUCTIVE BY 16.7%